OpenStudy (anonymous):

find the pattern 1,5,30,210,1680

OpenStudy (dape):

Well to go from the 1st to the 2nd you multiply by 5, from the 2nd to 3rd you multiply by 6 and so on.

OpenStudy (dape):

One way to write it mathematically is \[ a_1 = 1 \\ a_n = (n+3)a_{n-1} \] For example, the 4th number is: \[ a_4 = (4+3)a_{4-1} = 7a_3 = 7((3+3)a_{3-1}) = 42a_2 = 42((2+3)a_1) = 42*5*1 = 210 \]

OpenStudy (anonymous):

thanks